1. Steam at a temperature of 240 o C has a specific volume of 0.15436 m 3 /kg. Determine

pressure, specific internal energy, enthalpy and entropy.


1
Expert's answer
2021-05-05T07:40:47-0400
"p=1.48\ MPa\\u=2696\frac{kJ}{kg}\\h=2924.2\frac{kJ}{kg}\\s=6.717\frac{kJ}{kgK}\\"